An electron (m=9.11*10^ -31 kg,q=1.60*10^ -19 C) rest is accelerated through an electric potential difference of 12.0V. then enters a uniform magnetic field of 0.100 T (up), when initially moving (N)


a) Find the initial speed of the electron entering the magnetic field.


b) Find the magnetic force on the electron entering the magnetic field.


c) Find the radius of the circular path of the electron in the field .


1
Expert's answer
2021-03-29T09:02:07-0400

Mass of electron(m) = "9.11\times10^{-31}\space kg"

Charge of an electron(q) = "1.6\times10^{-19}\space C"

Magnetic Field(B) = "0.100 \space T"

Potential Difference(V) = "12\space V"


(a) Initial velocity of electron = "u\space" m/s

"\dfrac{1}{2}mu^2=eV"

"\Rightarrow u=\sqrt{\dfrac{2eV}{m}}"

"\Rightarrow u=2.06\times10^6\space m/s"


(b) "F=qvBsin\theta"

"\Rightarrow F=1.6\times10^{-19}\times2.06\times 10^6\times0.100\times sin90\degree"

"\Rightarrow F=3.296\times10^{-14}\space N"


(c) "r=mvqBsin\theta"

"\Rightarrow r=2.99\times10^{-44}\space m"